Aries:

Type the correct answer in each box. Use numerals instead of words. If necessary, use / for the fraction bar(s). Aiko had 17 comic books in her collection. She ordered some boxes of comic books. Each box contains 8 comic books. Aiko now has 65 comic books in her collection. Complete the equation that represents the total number of comic books in Aiko's collection, where b is the number of boxes ordered. ? b + ? = 65 b = ? Aiko receives some bonus boxes of comic books. She now has 97 comic books in her collection. Since Aiko received bonus boxes, the value of b has increased by ?

Arieonna:

well for the first question if it wants subtraction it would be 65-17=48 but if it was wanting addition it would be 65+17=82

KyledaGreat:

17 + 8b = 65 b = 6 Since Aiko received bonus boxes, the value of b has increased by an unknown amount, so we cannot determine the exact value without more information.

Aries:

Thank you so much!!! :)